100=5(2x+9) what is x
Found 2 solutions by maykim14, shree840:
Answer by maykim14(1)

x = 5.5
100=5(2x+9)
100/5=2x+9
20-9=2x
11/2=2x/2
x=5.5

Answer by shree840(260)

100=5(2x+9) open up the bracket . that means multiplying 5 with both 2x and 9.
100=10x+45 subtract 45 from both sides.
100-55=10x=45
x=4.5 ANS